# revtex4-2Repair the float

### v1

@@ IV Summary
We consider various kinematic distributions of the charged lepton. The models we present successfully reproduce the peak structure in data distributions.
Once these models are developed, they generate events significantly faster than “classical” generators. We also anticipate that these models can be adapted to more realistic scenarios after retraining them on experimental data. Essentially, they can serve as pre-trained models that can be fine-tuned for specific applications.
@@ Acknowledgements.
This research is partly (K.M.G., A.M.A., J.T.S.) or fully (B.E.K., J.L.B, H.P., R.D.B.) supported by the National Science Centre under grant UMO-2021/41/B/ST2/ 02778. K.M.G is partly supported by the “Excellence Initiative – Research University” for the years 2020-2026 at the University of Wrocław.

### v2

@@ IV Summary
Our model approach focuses on a simplified scenario in which the models, for neutrino energies ranging from $300$ MeV to $10$ GeV, predict muon kinematic variables—namely, its energy and scattering angle. We consider various kinematic distributions of the charged lepton (for neutrino energy from $300$ MeV to $10$ GeV). The models we present successfully reproduce the peak structure of the data distributions.
Once these models are developed, they generate events faster than “standard” generators [^5] Generating 1 million events on a single-core CPU takes approximately 12 minutes with NuWro (full simulation), and around 28 seconds with GAN. When GAN predictions are made on a typical gaming GPU (RTX 4080), the time is reduced to approximately 3 seconds.. We also anticipate that these models can be adapted to more realistic scenarios after retraining them on experimental data. Essentially, they can serve as pre-trained models that can be fine-tuned for specific applications.
@@ Acknowledgements.
This research is partly (K.M.G., A.M.A., J.T.S.) or fully (B.E.K., J.L.B, H.P., R.D.B.) supported by the National Science Centre under grant UMO-2021/41/B/ST2/ 02778. J.L.B. and J.T.S. are also supported by Polish Ministry of Science grant 2022/WK/15. K.M.G is partly supported by the “Excellence Initiative – Research University” for the years 2020-2026 at the University of Wrocław.
Appendix A GAN versus training data
In Fig. 16 we present comparison of the GAN predictions against training dataset. The training and GAN generated data are for neutrino energies larger than $950$ MeV and lower than $1050$ MeV.
Figure 16: Event histograms of the training subsamples and their GAN counterpart. Top row: QE histograms. Bottom row: INC histograms. Left column: $E_{\mu}$ distributions. Right column: $\theta$ distributions.